1452 mL of nitrogen is at 1.8 atm and 287 K.  What is the new temperature if the volume increases to 1500. mL and the pressure decreases to standard pressure?




1
Expert's answer
2021-02-10T01:32:21-0500

According to the combined gas law:

P1V1/T1 = P2V2/T2

where P1, V1 and T2 - initial pressure, volume and temperature, respectively; P2, V2, T2 - final pressure, volume and temperature, respectively.

From here:

T2 = P2V2T1 / P1V1 = 1 atm × 1500 mL × 287 K / (1.8 atm × 1452 mL) = 164.7 K


Answer: 164.7 K